OpenStudy (anonymous):

write the rational expression in lowest terms: x^2-6x+8 / x^2+2x-24

OpenStudy (anonymous):

\[\frac{x^2-6 x+8}{x^2+2 x-24}=\frac{((x-4) (x-2))}{((x-4) (x+6))}=\frac{x-2}{x+6} \]